I am having the toughest time figuring out this problem. x/3+y/5=-1, x-3y=-3. how do i solve this equation.

Found 2 solutions by mananth, Alan3354:
Answer by mananth(16949)

x/3+y/5=-1
lcd = 15
multiply by the LCD
5x+3y=-15...............1
x-3y=-3................2

multiply (1)by 1
Multiply (2) by 1
5x+3y=-15
1x-3y=-3

Add the two equations

6x=-18

/6

x=-3

plug value of x in (1)

5x+3y=-15
-15+3y= -15
3y=-15+15
3y=0
y=0


Answer by Alan3354(69443)

x/3+y/5=-1, x-3y=-3. how do i solve this equation.
It's 2 equations.
--------------
x/3+y/5=-1 eqn 1
x - 3y= -3 eqn 2
------------------
5x +3y = -15 eqn 1 x15
x - 3y = -3 eqn 2
-------------------- Add
6x = -18
x = -3
---------
etc